# PWDFT-SW: Extending the Limit of Plane-Wave DFT Calculations to 16K Atoms on the New Sunway Supercomputer

### v1

@@ IV-C Weak Scalability
Figure 14: Weak Scaling of PWDFT
The weak scaling of PWDFT for simulating systems consisting of 64 to 16,384 atoms is shown in Figure 14, from which we can see that PWDFT exhibits excellent weak scaling properties. Since the computational complexity of PWDFT is $O(N^{3})$, and the number of CGs is set to a quarter of atoms number, the ideal scaling should be $O(N^{2})$ which is represented by a dotted line in Figure 14.
@@
TABLE III: Comparision with Different PW-based DFT Software
@@
IV-D Speedup Compared with Master Version
Figure 15: Speedup for Si4096 with master version
The results show that when only adopting CPE to accelerate, the speedup varies from 100.0 to 7.9, and decreases dramatically with the increase of the number of CGs. For comparison, when utilizing all the mechanisms proposed in Section III, the acceleration effect is noticeably enhanced. Under more common circumstances, the fully optimized version achieves 64.8 faster than the master version when we adopt one-quarter of the atom number, i.e. 1,024 CGs.
TABLE IV: Minimun CGs required to calculate the system
System Master Version Optimized Version
Si1000 512 25
Si4096 512 64
Si8192 out of memory 512
G11520 out of memory 720
G16384 out of memory 2048
Meanwhile, memory optimization in Section III-A and Sectioin III-B significantly extends the limit of the studied system for PWDFT, As shown in Table  IV, The master version requires at least 512 CGs to calculate a silicon system containing 4,096 atoms (Si4096). However, after our optimization, only 64 CGs are needed. Furthermore, only 512 CGs are required for a silicon system with 8,192 atoms (Si8192), which is quite impressive for PW-based DFT calculations. Consequently, the limit of PWDFT is extended from a physical system of 4,096 atoms to one encompassing 16,384 atoms.
V Discussion
In terms of computational ability, PWDFT is still memory-bound for the simulation of larger systems on the new Sunway supercomputer. Memory bottleneck occurs in the storage of pseudopotential and Hamiltonian matrices. To our knowledge, our work already reached the boundary of PW-based DFT on the new Sunway supercomputer. To further expand the capabilities of PW-based DFT software on the new Sunway supercomputer, one must either (1) create novel algorithms with reduced memory complexity, or (2) develop numerical libraries that align with the large memory scheme of the new Sunway supercomputer. It is important to emphasize that this work does not primarily focus on these two aspects.
@@ VI-A Other Parallel PW-based Software
There are several software packages based on plane wave available for performing DFT, such as Vienna Ab initio Simulation Package (VASP), Qbox, Quantum ESPRESSO (QE), PEtot [24] and PWDFT [19]. VASP, Qbox, and QE are traditional software mainly used for simulations of medium-size atoms (hundreds of atoms) and PWDFT and PEtot are designed for large-size atoms (thousands of atoms). The more detailed comparison is shown in table  III.
@@
Recently years, thanks to heterogeneous computing systems such as graphics process units (GPU) and many-core microarchitecture, those traditional software packages come to rejuvenated after being ported onto the new generation of supercomputers. In 2006, the winner of the Gordon Bell Prize extended the limit of Qbox on the BlueGene/L supercomputer to 1K atoms, scaling up to 128k cores on 65,536 nodes [13]. In 2017, researchers successfully ran Qbox [7] on 9418 nodes on Trinity with 8,800 atoms. In 2013, Jia [24] implemented PEtot on a multi-node GPU machine with x13-x22 speedups over the CPU calculations for a typical 512-atom system. Researchers ported VASP to Sunway TaihuLight [39] and achieved 2.90x to 4.48x speedup, but they only focused on small cases containing 32 Ge atoms and one Ru atom. It is significantly smaller than other works.
@@ VI-B Other Research Works on the New Sunway Supercomputer
There are a lot of successful works that effectively utilize the massive computing resource of the new Sunway supercomputer [18, 40, 34, 36], for instance: (1) By implementing a novel sparse direct solver PEXSI, DGDFT extended the limit of adaptive local basis DFT calculation to 2.5 million atoms using 35.9 million cores, achieve  5% theoretical peak performance [18]. (2) A brand new computational framework to solve quantum many-body problems by Artificial Intelligence (AI) was proposed and adapted to the Sunway, scaling up to 40 million heterogeneous cores with high efficiency [40]. (3) Based on the abstraction of application and hierarchical hardware, researchers presented a unified programming model, UniQ, for multiple simulation methods. UniQ can generate execution with high performance without any human involvement. It scales to 399,360 nodes and achieves up to 28.59x speedup over state-of-art framework [38].
@@ VII Conclusion
In summary, we have significantly optimized the ability and performance of PWDFT on the new Sunway supercomputer. Firstly, we reduce memory usage and improve memory bandwidth utilization by refactoring pseudopotential calculations and in-place data transforming, so that a single CG with 16 GB memory is able to hold much more wavefunctions and a larger Hamiltion matrix for larger systems. Secondly, we accelerate PWDFT by adopting CPEs, and granularity-aware parallelization to take full advantage of physical floating-point performance. Meanwhile, we also present multistage Allreduce to to optimize network bandwidth utilization. As a result, PWDFT can now perform calculations of 4096 atoms using 64 CGs, 720 CGs for 11,520 atoms and 2,048 CGs for 16,384 atoms. These optimizations lead to a 64.8x speedup when simulating a silicon system comprising 4,096 atoms and utilizing 1,024 CGs. Additionally, our experiments demonstrate that our parallel design can achieve good strong scalability, even when utilizing a large number of cores.

### v2

@@ IV-C Weak Scalability
Figure 14: Weak Scaling of PWDFT w.r.t. the number of electrons.
The weak scaling of PWDFT for simulating systems consisting of 64 to 16,384 atoms is shown in Figure 14, from which we can see that PWDFT exhibits excellent weak scaling properties. Since the computational complexity of PWDFT is $O(N^{3})$, and the number of CGs is set to a quarter of atoms number, the ideal scaling should be $O(N^{2})$, which is represented by a dotted line in Figure 14.
@@
IV-D Speedup Compared with Master Version
Figure 15: Speedup for Si4096 with master version.
The results show that when only adopting CPE to accelerate, the speedup varies from 100.0 to 7.9, and decreases dramatically with the increase of the number of CGs. For comparison, when utilizing all the mechanisms proposed in Section III, the acceleration effect is noticeably enhanced. Under more common circumstances, the fully optimized version achieves 64.8 times faster than the master version when we adopt one-quarter of the atom number, i.e., 1,024 CGs.
TABLE III: Minimum CGs required to calculate the system.
System Master Version Optimized Version
Si1000 512 25
Si4096 512 64
Si8192 Out of Memory 512
G11520 Out of Memory 720
G16384 Out of Memory 2048
Meanwhile, memory optimization in Section III-A and Section III-B significantly extends the limit of the studied system for PWDFT. As shown in Table III, The master version requires at least 512 CGs to calculate a silicon system containing 4,096 atoms (Si4096). However, after our optimization, only 64 CGs are needed. Furthermore, only 512 CGs are required for a silicon system with 8,192 atoms (Si8192), which is quite impressive for PW-based DFT calculations. Consequently, the limit of PWDFT is extended from a physical system of 4,096 atoms to one encompassing 16,384 atoms.
IV-E Speedup with Different Energy Cutoff
To comprehensively examine the performance of our optimization strategies across various computational conditions, we conducted additional tests by varying the plane-wave energy cutoff ($E_{cut}$). We selected four different $E_{cut}$ values (5, 10, 15, and 20 Ha), representing both our default setting (5 Ha) and more realistic values typically encountered in practical DFT calculations. We compared the SCF iteration times of three versions of our implementation: the Master version (MPE, utilizing only the main processing element), the version with additional computing processing elements (MPE+CPE, including slave cores), and the fully optimized version (All OPT), incorporating all optimization techniques proposed in this work. The results are summarized in Figure 16.
Figure 16: SCF iteration time and speedup comparisons at different $E_{cut}$ values.
From Figure 16, we observe the following:
  • At all tested energy cutoffs, both the MPE+CPE and All OPT versions significantly outperform the Master version, clearly demonstrating that utilizing slave cores effectively accelerates SCF calculations.
  • The speedup of the MPE+CPE version relative to the Master version improves as the energy cutoff increases, achieving approximately 2.7$\times$, 5.2$\times$, 6.8$\times$, and 9.5$\times$ at 5, 10, 15, and 20 Ha, respectively. This trend underscores the effectiveness of parallelization using additional computing elements as computational intensity increases.
  • More importantly, the fully optimized All OPT version, which includes all optimizations discussed in this paper, consistently delivers superior performance. Specifically, it achieves a speedup of approximately 3.5$\times$ at 5 Ha, which steadily rises to approximately 7.8$\times$, 9.5$\times$, and 11.8$\times$ at 10, 15, and 20 Ha, respectively. This substantial improvement further validates the robustness and practical applicability of our proposed optimization strategies.
These results confirm that the optimization methods proposed in this study are effective across a broad range of energy cutoffs, ensuring that our conclusions remain valid and relevant under realistic computational scenarios commonly encountered in large-scale PW-based DFT simulations.
V Discussion
In terms of computational ability, PWDFT is still memory-bound for the simulation of larger systems on the new Sunway supercomputer. Memory bottleneck occurs in the storage of pseudopotential and Hamiltonian matrices. To our knowledge, our work has already reached the boundary of PW-based DFT on the new Sunway supercomputer. To further expand the capabilities of PW-based DFT software on the new Sunway supercomputer, one must either (1) create novel algorithms with reduced memory complexity or (2) develop numerical libraries that align with the large memory scheme of the new Sunway supercomputer. It is important to emphasize that this work does not primarily focus on these two aspects.
V-A Potential Applications of PWDFT-SW
The optimization techniques presented in this work are primarily developed and evaluated for the Sunway supercomputer; nevertheless, many of these strategies are not inherently limited to the Sunway architecture and can be generalized to other state-of-the-art supercomputing platforms, including GPUs [41, 42], Fugaku [45], and Summit [52].
For instance, the memory accessing optimizations and pseudopotential refactoring schemes, designed to reduce redundant communication and enhance data locality, are directly applicable to GPU-based platforms such as Summit, which benefit significantly from efficient memory management and reduced communication overhead. Similarly, the granularity-aware parallelization method can be adapted for other many-core CPU architectures, such as Fugaku, where appropriate resource allocation and optimized communication patterns can greatly enhance computational performance.
Future work will involve porting and tailoring these optimization strategies explicitly for GPU-based and ARM-based supercomputers. Such efforts would enable detailed benchmarking across different architectures, highlighting the universality and adaptability of our proposed optimizations. Additionally, further investigation into platform-specific tuning and potential new bottlenecks on these architectures would provide valuable insights, thereby broadening the applicability and impact of our methods within the high-performance computing community.
TABLE IV: List of Different PW-based DFT Software.
@@ V-A Potential Applications of PWDFT-SW
V-B Integration with Machine Learning Methods
Machine learning (ML) techniques have demonstrated significant potential in accelerating and enhancing DFT calculations. Recent prominent examples include DeepH [36, 51], which employs graph-based neural networks to predict DFT Hamiltonians directly from atomic structures, thereby circumventing expensive self-consistent iterations. Another noteworthy development is DeepKS [7, 8], which uses neural-network corrections to improve the accuracy of conventional DFT exchange-correlation functionals by learning from high-accuracy reference data. Additionally, DeepMD [48, 56] has revolutionized ab initio molecular dynamics simulations by using ML-driven interatomic potentials, significantly enhancing both computational efficiency and scalability. The remarkable success of DeepMD, highlighted by the ACM Gordon Bell Prize [37], underscores the transformative impact of ML in enabling large-scale molecular dynamics simulations with DFT-level accuracy. Furthermore, recent generative models such as MatterGen [57] have also demonstrated promising capability in efficiently generating stable inorganic materials, potentially complementing ML-accelerated DFT calculations. Additionally, the recent N2AMD framework [58], employing E(3)-equivariant neural networks to predict DFT Hamiltonians for nonadiabatic molecular dynamics, offers a promising direction for further accelerating PWDFT-based simulations of large-scale materials.
Beyond these specific frameworks, general ML strategies like Behler-Parrinello neural networks [17] and Gaussian Approximation Potentials (GAP) [33] have also been integrated successfully with DFT to create computationally efficient surrogate models, reducing computational demands while preserving high fidelity. Integrating such ML-based methods with the optimization strategies proposed in this study holds substantial promise. Future investigations could involve combining our parallel optimization methods with ML-based approximations to further enhance efficiency, particularly in large-scale materials simulations. Such integration could expand the application of PWDFT to even larger and more complex systems, offering avenues for significant future research.
@@ VI-A Other Parallel PW-based Software
There are several software packages based on plane wave available for performing DFT, such as Vienna Ab initio Simulation Package (VASP), Qbox, Quantum ESPRESSO (QE), PEtot [29], and PWDFT [24]. VASP, Qbox, and QE are traditional software mainly used for simulations of medium-size atoms (hundreds of atoms), and PWDFT and PEtot are designed for large-size atoms (thousands of atoms). We list the related works in Table IV. It is important to note that the list provided in Table IV involves results obtained from different physical systems and computational resources, thus serving primarily as a general reference rather than a rigorous benchmarking. The current work emphasizes performance optimization tailored specifically for the Sunway architecture, highlighting scalability and capability for handling large-scale systems. A detailed comparison under consistent accuracy settings and computational conditions has been reported separately in PyPWDFT [14].
@@
Recently, thanks to heterogeneous computing systems such as graphics processing units (GPU) and many-core microarchitectures, those traditional software packages have come to rejuvenate after being ported onto the new generation of supercomputers. In 2006, the winner of the Gordon Bell Prize extended the limit of Qbox on the BlueGene/L supercomputer to 1K atoms, scaling up to 128k cores on 65,536 nodes [18]. In 2017, researchers successfully ran Qbox [10] on 9418 nodes on Trinity with 8,800 atoms. In 2013, Jia [29] implemented PEtot on a multi-node GPU machine with x13-x22 speedups over the CPU calculations for a typical 512-atom system. Researchers ported VASP to Sunway TaihuLight [60] and achieved 2.90x to 4.48x speedup, but they only focused on small cases containing 32 Ge atoms and one Ru atom. It is significantly smaller than other works.
@@ VI-B Other Research Works on the New Sunway Supercomputer
There are a lot of successful works that effectively utilize the massive computing resource of the new Sunway supercomputer [23, 61, 47, 50], for instance: (1) By implementing a novel sparse direct solver PEXSI, DGDFT extended the limit of adaptive local basis DFT calculation to 2.5 million atoms using 35.9 million cores, achieving  5% theoretical peak performance [23]. (2) A brand new computational framework to solve quantum many-body problems by Artificial Intelligence (AI) was proposed and adapted to the Sunway, scaling up to 40 million heterogeneous cores with high efficiency [61]. (3) Based on the abstraction of application and hierarchical hardware, researchers presented a unified programming model, UniQ, for multiple simulation methods. UniQ can generate execution with high performance without any human involvement. It scales to 399,360 nodes and achieves up to 28.59x speedup over the state-of-the-art framework [59].
@@ VII Conclusion
In summary, we have significantly optimized the ability and performance of PWDFT on the new Sunway supercomputer. Firstly, we reduce memory usage and improve memory bandwidth utilization by refactoring pseudopotential calculations and in-place data transforming, so that a single CG with 16 GB memory is able to hold much more wavefunctions and a larger Hamiltonian matrix for larger systems. Secondly, we accelerate PWDFT by adopting CPEs and granularity-aware parallelization to take full advantage of physical floating-point performance. Meanwhile, we also present a multistage Allreduce to optimize network bandwidth utilization. As a result, PWDFT can now perform calculations of 4096 atoms using 64 CGs, 720 CGs for 11,520 atoms, and 2,048 CGs for 16,384 atoms. These optimizations lead to a 64.8x speedup when simulating a silicon system comprising 4,096 atoms and utilizing 1,024 CGs. Additionally, our experiments demonstrate that our parallel design can achieve good strong scalability, even when utilizing a large number of cores.
Acknowledgment
We thank the reviewers of ICPP 2023, IPDPS 2024, Euro-Par 2024, and TPDS for their feedback to strengthen this work. This work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences (XDB0500102, XDB0450101, XDB1170000), the Laoshan Laboratory (LSKJ202300305), the Innovation Program for Quantum Science and Technology (2021ZD0303306), the National Natural Science Foundation of China (22173093, 22373096, 21688102), the Anhui Provincial Key Research and Development Program (2022a05020052), the National Key Research and Development Program of China (2021YFB0300600), the CAS Project for Young Scientists in Basic Research (YSBR-005), and the Fundamental Research Funds for the Central Universities (WK2320000061).